OpenStudy (anonymous):

very urgent the market price is 10% higher than cost price.a shopkeeper offers 10% discount on market price. does it gain or loss

OpenStudy (anonymous):

loss

OpenStudy (anonymous):

how

OpenStudy (anonymous):

for example. if the cost of other market is 100, this market makes 10% higher, that means the cost is 110. and discount 10% of 110 is discount $11; 110-11=99. the 10% discount based on the new price , not the original price. the new price as high, the discount money as high. right?

OpenStudy (anonymous):

don't be confused by 10% high and then10% discount as the same. after making higher price, then discount from that high means substract higher. do you got it?

OpenStudy (anonymous):

if the discount is 8%, it's another topic. they gain

OpenStudy (anonymous):

everybody gives out the idea , please. don't just watching me like that. i'm gonne

OpenStudy (anonymous):

well can u solve this sum or not

OpenStudy (anonymous):

what do you mean? I just give the example. if you want to solve for it, you must have the price, how can i solve without the price?

OpenStudy (anonymous):

okay can we assume that price be x or just see my solution vat i did CP = x MP= 11x/10 Discount = 10% SP = 9/10 * 11x/10 SP = 99x/100 CP - SP = LOSS x - 99x/100 x/100 Loss% = x/100 / x * 100 = 1 % loss

OpenStudy (anonymous):

perfect! you are amazing

OpenStudy (anonymous):

I don't know whether it's right or wrong, I cannot do that. that's what i mean. I solve any problems in simplest way I think. put everything in logic like that is tooooo far from mine

OpenStudy (anonymous):

k